detemine x so that 800=8*10^8*x^-3/2
please give me the answer as soon as possible its answer is is 10^4
i m very thankful to u

Answer by stanbon(75887)

800=8*10^8*x^-3/2
divide both sides by 8 to get:
100 = 10^8 x^(-3/2)
Multiply both sides by x^(3/2)
100 x^(3/2) = 10^8
x^(3/2) = 10^6
Raise both sides to the (2/3) power
x = 10^4
x = 10,000
==============
Cheers,
Stan H.